OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −one halfx − 2 y = 2x − 2 y = −one halfx − 7

OpenStudy (anonymous):

@RadEn ?

OpenStudy (anonymous):

@jhonyy9 ?

OpenStudy (anonymous):

@Rogue

OpenStudy (rogue):

Hiya.

OpenStudy (rogue):

The slope intercept form of a line is \[y = mx + b\] where m is the slope and b is the y-intercept. In order for lines to be parallel, they have to have the same slopes, so which answer choices can you eliminate?

OpenStudy (anonymous):

I could eliminate the first and last choices?

OpenStudy (rogue):

Well the slope of the line y = 2x + 2 is 2, so the slopes of your parallel lines will also have to be 2. The lines in choice 1 & 3 have slopes 2, so they can possibly be the answer. The lines in choice 2 & 4 do not have slope 2, so they can't be parallel and so you can eliminate them.

OpenStudy (rogue):

So out of choice 1 & 3, Choice 1: y = 2x − 7 Choice 3: y = 2x − 2 Which one passes through the point (5, 3)?

OpenStudy (anonymous):

Choice 1.? Sorry, I'm horrible at math.

OpenStudy (rogue):

Yeah, it's 1, good job :) When you plug in the point (5, 3) into the equation, you should get both sides to be equal. y = 2x − 7 <--- (5, 3} 3 = 2*5 - 7 3 = 10 - 7 3 = 3 That makes sense and so the point (5, 3) is on that line.

OpenStudy (anonymous):

Awesome. :) Thanks so much.

OpenStudy (rogue):

No problem =)